# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

### v1

@@ 4.1.1. Ensemble learning
Ensemble learning in deep static MaE recognition can exploit complementary information from multiple feature representations of a single emotion image. More specifically, ensemble learning can be used at different stages of the deep static MaE recognition, such as, data preprocessing (Kim et al., 2016), input enhancement (Sharifnejad et al., 2021; Hariri and Farah, 2021; Mohan et al., 2021; Shao and Qian, 2019; Viswanatha Reddy et al., 2020), network generation (Kim et al., 2015; Pons and Masip, 2018b), feature extraction (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), and emotion classification (Mohan et al., 2021; Hua et al., 2019). More specifically, deformation and normalization have been used to preprocess the data before training the deep models for MaE recognition (Kim et al., 2016). When applied at the stage of input enhancement, ensemble learning can improve the MaE recognition by integrating multiple types of textural features, such as, local binary patterns (LBPs) (Sharifnejad et al., 2021; Hariri and Farah, 2021), facial landmark point (Mohan et al., 2021), covariance feature (Shao and Qian, 2019), gradient images (Hariri and Farah, 2021), and gravitational force (Viswanatha Reddy et al., 2020). At the stage of network generation, deep models with various kernel shapes and parameter initializations are ensembled to improve performance (Kim et al., 2015; Pons and Masip, 2018b). For example, Saurav et al. (Saurav et al., 2022) proposed an integrated convolutional neural network (CNN) architecture that leveraged two deep models with distinct kernel shapes. The aggregation of features extracted from deep models represents another prevalent research direction for constructing model ensembles in deep static MaE recognition. Several studies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023) have explored integrating different regions (e.g., eyes, mouth, nose, and the entire image) of the face for model ensembling. Through the different ensemble strategies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), an integrated feature that combines information from all facial regions can be generated for MaE recognition. In addition to the common integration rules (e.g., majority voting, simple average, and weighted average), the weighted average for decision ensemble was also investigated in (Mohan et al., 2021; Hua et al., 2019). Despite the aforementioned merits, ensemble learning increases the training cost of the model and lacks transparency and interoperability.
@@ 4.1.2. Transfer learning
Transfer learning in deep static MaE recognition enables deep models to be fine-tuned on MaE datasets after being pre-trained on large-scale, high-quality MaE datasets. The utilization of transfer learning can alleviate the overfitting issue due to the limited training data for static MaE recognition. By pre-training the model with additional data from other relative tasks, deep static MaE recognition methods can obtain more generic knowledge. For example, Ng et al. (Ng et al., 2015) used ImageNet (Deng et al., 2009) dataset to obtain the pre-trained models that are fine-tuned by the other related MaE datasets. The training efficiency can be improved by only fine-tuning the dense layers of deep models; therefore, the training computing expenditure is reduced in (Akhand et al., 2021). Instead of using general-purpose datasets, MaE-specific datasets can be used for deep model pre-training (Ngo and Yoon, 2020; Atabansi et al., 2021). For instance, Ngo et al. (Ngo and Yoon, 2020) employed the face identification dataset (e.g., VGGFace2 (Cao et al., 2018)) for model transfer. To further mitigate the overfitting in deep static MaE recognition, Atabansi et al. (Atabansi et al., 2021) used another MaE recognition dataset with high resolution and huge quantity for model pre-training.
@@ 4.1.3. Multi-task learning
Multi-task learning allows a deep model to extract macro-expression features from static facial images by using other facial behavior analysis tasks as auxiliary tasks in the deep static MaE recognition. For example, recent deep static MaE recognition methods have integrated facial landmark localization (Chen et al., 2021) and AU detection (Pons and Masip, 2018a) to extract more robust MaE features (Pons and Masip, 2018a; Chen et al., 2021; Yu and Xu, 2022; Liu et al., 2024). Moreover, Pons and Masip found that leveraging shared features across AU and MaE can enhance the performance of recognition (Pons and Masip, 2018a). To further improve the synergy, Li et al. (Li et al., 2023) introduced an alignment loss to constrain the feature distribution between the AU detection and MaE recognition tasks. Furthermore, the multi-task learning can be used to fuse the global and local facial expressions by automatically assigning weights based on the importance of global and local facial information (Yu et al., 2020a; Xiao et al., 2023). Besides, facial expression synthesis (Zhao et al., 2021b), head pose estimation (Chen et al., 2022b), body gesture (Zaghbani and Bouhlel, 2022), and gender learning (Pan et al., 2022) had been demonstrated as promising collaborative auxiliary tasks for deep static MaE recognition and could significantly improve performance. Therefore, Chen et al. (Chen et al., 2022b) proposed a dual-attention based multi-task learning framework that contains a separate channel attention mechanism to calculate task-specific attention weights and an orthogonal channel attention loss to optimize the selection of feature channels for each auxiliary task. Since human emotion is conveyed equally via the body and the face in most cases, Zaghbani et al. (Zaghbani and Bouhlel, 2022) combined the upper body gesture action detection task for the deep static MaE recognition. Pan et al. (Pan et al., 2022) incorporated gender learning as an auxiliary task to factor in the effects of gender on the deep static MaE recognition since the characteristics of the same facial expressions from males, females, and infants can be significantly different. Instead of using a single auxiliary task, several studies proposed various multiple tasks for deep static MaE recognition to achieve more synergy. For instance, Qin et al. (Qin et al., 2024) developed an algorithm that simultaneously performs face recognition, MaE recognition, age estimation, and face attribute prediction. Similarly, Foggia et al. (Foggia et al., 2023) proposed a comprehensive multi-task framework to integrate gender, age, ethnicity, and MaE recognition using facial images.
@@ 4.1.4. Attention-based learning
Attention-based learning is used to enrich the spatial information in deep static MaE recognition. The attention modules can focus on both global and local regions and generate more comprehensive static facial expression features that typically include local information, global information, and the corresponding dependency. Drawing inspiration from the human ability to locate salient objects in complex visual environments, attention mechanisms have emerged as a powerful tool in deep static MaE recognition (Itti and Koch, 2001). By prioritizing critical facial regions, attention modules can significantly enhance the extraction and integration of expression features in order to achieve a more accurate and robust recognition model. For example, Fernandez et al. (Marrero Fernandez et al., 2019) developed an attention module to enhance the extraction of expression features by assigning higher weights to relevant regions. Li et al. (Li et al., 2020a) integrated LBP and deep features with an attention mechanism to enhance the performance of deep static MaE recognition. In wild scenarios, pose variations and occlusions present significant challenges for deep static MaE recognition (Li et al., 2019b; Wang et al., 2020b). To address the aforementioned issues, various attention modules have been proposed to enhance model robustness and performance. For example, Li et al. (Li et al., 2019b) introduced an attention module designed to identify local facial regions associated with MaEs while simultaneously incorporating complementary global facial information. This dual-focus approach enables more robust MaE recognition under occlusion conditions by balancing localized details with a holistic view of the face. Wang et al. (Wang et al., 2020b) proposed a region attention module that is tailored to address occlusions and pose variations. The region attention module in (Wang et al., 2020b) highlights significant facial regions by leveraging attention weights and combines local region features to generate a compact and fixed-length feature representation for final classification. Motivated by the idea of integrating global and local features, Zhao et al. (Zhao et al., 2021a) developed a global multi-scale and local attention module. Zhao et al. (Zhao et al., 2021a) confirmed that the proposed attention module can ensure a comprehensive representation of facial expressions by capturing global features through multiple receptive fields while guiding the model to focus on salient local features. Motivated by facial attributes and human perception mechanisms, Liu et al. (Liu et al., 2022a) introduced an adaptive multilayer perceptual attention module. By effectively capturing critical information from global, local, and salient facial regions, the proposed module in (Liu et al., 2022a) significantly improves the robustness of deep static MaE recognition.
@@
First proposed in (Dosovitskiy et al., 2020), vision transformers (ViTs) have brought remarkable advancements in the domain of deep static MaE recognition, primarily due to their ability to capture long-range dependencies across image patches (Aouayeb et al., 2021; Zheng et al., 2023). By leveraging the self-attention module inherent in transformers, Aouayeb et al. (Aouayeb et al., 2021) pioneered the integration of ViTs with a squeeze-and-excitation block that can extract expression-related features. Xue et al. (Xue et al., 2021) exploited the ViT architecture to explore relationships among various facial regions for a more holistic understanding of static MaEs. To address the occlusion issue, Liu et al. (Liu et al., 2023a) introduced a patch attention module that assigns attention weights to local facial regions. The integration of the patch attention module with a ViT can effectively capture both local and global dependencies. Building upon the advancements of ViTs, Zheng et al. (Zheng et al., 2023) developed a dual-stream model that leverages a cross-fusion mechanism to combine facial landmark features with image-based features. Furthermore, the incorporation of additional modalities has proven instrumental in augmenting transformer-based deep static MaE recognition. For example, gradient images (Meng and Shi, 2024) and gray-level co-occurrence matrices (Sun et al., 2023a) have been used as complementary inputs to enrich the feature space. These modalities not only enhance the discriminative power of the models but also provide deeper insights into the complex interplay of facial features. The optimization of transformer architectures also plays a critical role in advancing deep static MaE recognition. In (Feng et al., 2023), Feng et al. demonstrated that adjusting the parameter structures of ViTs through optimization algorithms has led to significant performance improvements.
@@ 4.1.5. Self-supervised learning
The self-supervised learning had also been applied to multi-modal MaE recognition to address the challenges of integrating data from various modalities without explicit annotation. Siriwardhana et al. (Siriwardhana et al., 2020) utilized pre-trained self-supervised models to extract text, speech, and vision features to improve deep static MaE recognition tasks. Moreover, multi-modal self-supervised learning frameworks had been explored to extract and fuse multi-modal features to achieve promising results in deep static MaE recognition (Chaudhari et al., 2023; Halawa et al., 2024).
4.1.6. Potential limitations of different methods on deep static MaE recognition.
While various methods in deep static MaE recognition have significantly improved model performance and broadened their applicability, they still present certain limitations. Ensemble learning, which combines multiple models to boost performance, often increases computational costs and reduces transparency and interoperability due to added complexity. Transfer learning, which relies on pre-trained models, can be less effective if there is a significant domain gap between the source and target datasets. Moreover, improper fine-tuning can reduce the model’s ability to adapt to new datasets. Multi-task learning enhances model performance by jointly learning MaE recognition alongside related tasks. However, the choice of auxiliary tasks is critical since irrelevant or conflicting tasks may hinder learning. Attention mechanisms, while powerful, can focus excessively on certain regions, potentially ignoring other important information, especially in cases of facial occlusions or pose variations. Self-supervised learning is very dependent on the quality of input data or the limited labeled data. Most current large-scale MaE datasets are collected from different viewing angles, lighting conditions, and occlusions in network environments that cannot be controlled. Therefore, if the unlabeled data contains much noise or atypical samples, then the model may learn incorrect feature representations, which impairs its final recognition performance.
@@
Explicit Spatio-temporal Learning Zhang et al. (Zhang et al., 2017) 2017 DA N/A CK+/Oulu-CASIA/MMI 45.5%/25.8%/24.4% CNN+RNN 10 fold
@@
Hasani et al. (Hasani and Mahoor, 2017) 2017 FA+DA RES CK+/MMI/FERA/DISFA 67.52%/54.76%/41.93%/40.51% CNN+RNN 5 fold
@@
Multi-task Learning Yu et al. (Yu et al., 2020b) 2020 FA+FD N/A CK+/Oulu-CASIA 98.77/90.40% CNN+LSTM 10 fold
Jin et al. (Jin et al., 2021) 2021 DA RES Aff-wild2 89.87% CNN Hold out
Xie et al. (Xie et al., 2023) 2023 FA+DA+ROI Trans+Atten MU-MOSI/CMU-MOSEI 84.61%/84.61% BERT Hold out
Attention-Based Learning Meng et al. (Meng et al., 2019) 2019 FD+FA RES+Atten CK+/AFEW 8.0 99.69%/51.8% CNN 10 fold
@@
Huang, et al. (Huang et al., 2021) 2021 FA RES+Atten CK+/FERplus/RAF-DB 100%/90.04%/88.26% CNN 10 fold+Hold out
@@
Chen et al. (Chen et al., 2024) 2024 DA RES+Trans DFEW/FERV39k/AFEW 8.0
@@
Zhang et al. (Zhang et al., 2022) 2023 FD+DA Trans MAFW11/MAFW43/DFEW/AFEW
@@
Sun et al. (Sun et al., 2023b) 2023 ROI Trans DFEW/FERV39k/MAFW
@@
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; FA: Face alignment; N/A: Not applicable;
  • 2Trans: Transformer; Atten: Attention; nsCNN: n-stream CNN; LOSO: Leave-one-subject-out; RNN: Recurrent neural network; LSTM: Long short-term memory; BERT: Bidirectional encoder representations from transformer; GCN: Graph convolutional network;
  • 3 UAR: Mean class accuracy; WAR: Overall accuracy; RMSE: Root mean squared error; CC: Correlation coefficient; CCC: Concordance correlation coefficient;

### v3

@@ 4.1.1. Ensemble learning
Ensemble learning in deep static MaE recognition can exploit complementary information from multiple feature representations of a single emotion image. Ensemble learning can be used at different stages of the deep static MaE recognition, such as, data preprocessing (Kim et al., 2016), input enhancement (Sharifnejad et al., 2021; Hariri and Farah, 2021; Mohan et al., 2021; Shao and Qian, 2019; Viswanatha Reddy et al., 2020), network generation (Kim et al., 2015; Pons and Masip, 2018a), feature extraction (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), and emotion classification (Mohan et al., 2021; Hua et al., 2019). More specifically, deformation and normalization have been used to preprocess the data before training the deep models for MaE recognition (Kim et al., 2016). When applied at the stage of input enhancement, ensemble learning can improve the MaE recognition by integrating multiple types of textural features, such as, local binary patterns (LBPs) (Sharifnejad et al., 2021; Hariri and Farah, 2021), facial landmark point (Mohan et al., 2021), covariance feature (Shao and Qian, 2019), gradient images (Hariri and Farah, 2021), and gravitational force (Viswanatha Reddy et al., 2020). At the network generation stage, deep models with various kernel shapes and parameter initializations are ensembled to improve performance (Kim et al., 2015; Pons and Masip, 2018a). For example, Saurav et al. (Saurav et al., 2022) proposed an integrated convolutional neural network (CNN) architecture that leveraged two deep models with distinct kernel shapes. The aggregation of features extracted from deep models represents another prevalent research direction for constructing model ensembles in deep static MaE recognition. Several studies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023) have explored integrating different regions (e.g., eyes, mouth, nose, and the entire image) of the face for model ensembling. Through the different ensemble strategies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), an integrated feature can combine all facial regions for MaE recognition. Besides the common integration rules (e.g., majority voting, simple average, and weighted average), the weighted average for decision ensemble was also investigated in (Mohan et al., 2021; Hua et al., 2019). Despite the aforementioned merits, ensemble learning increases the training cost and lacks transparency and interoperability.
@@ 4.1.2. Transfer learning
Transfer learning in deep static MaE recognition enables deep models to be fine-tuned on MaE datasets after being pre-trained on large-scale, high-quality related datasets. The utilization of transfer learning can alleviate the overfitting issue due to the limited training data for static MaE recognition. By pre-training the model with additional data from other relative tasks, deep static MaE recognition methods can obtain more generic knowledge. For example, Ng et al. (Ng et al., 2015) used ImageNet (Deng et al., 2009) dataset to obtain the pre-trained models that are fine-tuned by the other related MaE datasets. To further enhance training efficiency, (Akhand et al., 2021) fine-tuned only the dense layers of deep models, thereby reducing computational costs. Instead of using general-purpose datasets, MaE-specific datasets can be used for deep model pre-training (Ngo and Yoon, 2020; Atabansi et al., 2021). For instance, Ngo et al. (Ngo and Yoon, 2020) employed the face identification dataset (e.g., VGGFace2 (Cao et al., 2018)) for model transfer. To further mitigate the overfitting in deep static MaE recognition, Atabansi et al. (Atabansi et al., 2021) used another MaE recognition dataset with high resolution and huge quantity for model pre-training.
@@ 4.1.3. Multi-task learning
Multi-task learning allows a deep model to extract macro-expression features from static facial images by using other facial behavior analysis tasks as auxiliary tasks in the deep static MaE recognition. For example, recent deep static MaE recognition methods have integrated facial landmark localization (Chen et al., 2021) and AU detection (Pons and Masip, 2018b) to extract more robust MaE features (Pons and Masip, 2018b; Chen et al., 2021; Yu and Xu, 2022; Liu et al., 2024). Moreover, Pons and Masip found that leveraging shared features across AU and MaE can enhance the performance of recognition (Pons and Masip, 2018b). To further improve the synergy, Li et al. (Li et al., 2023) introduced an alignment loss to constrain the feature distribution between the AU detection and MaE recognition tasks. Furthermore, the multi-task learning can be used to fuse the global and local facial expressions by automatically assigning weights based on the importance of global and local facial information (Yu et al., 2020; Xiao et al., 2023). Besides, facial expression synthesis (Zhao et al., 2021a), head pose estimation (Chen et al., 2022a), body gesture (Zaghbani and Bouhlel, 2022), and gender learning (Pan et al., 2022) had been demonstrated as promising collaborative auxiliary tasks for deep static MaE recognition and could significantly improve performance. Therefore, Chen et al. (Chen et al., 2022a) proposed a dual-attention based multi-task learning framework that contains a separate channel attention mechanism to calculate task-specific attention weights and an orthogonal channel attention loss to optimize the selection of feature channels for each auxiliary task. Since human emotion is conveyed equally via the body and the face in most cases, Zaghbani et al. (Zaghbani and Bouhlel, 2022) combined the upper body gesture action detection task for the deep static MaE recognition. Pan et al. (Pan et al., 2022) incorporated gender learning as an auxiliary task to factor in the effects of gender on the deep static MaE recognition since the characteristics of the same facial expressions from males, females, and infants can be significantly different. Instead of using a single auxiliary task, several studies proposed various multiple tasks for deep static MaE recognition to achieve more synergy. For instance, Qin et al. (Qin et al., 2024) developed an algorithm that simultaneously performs face recognition, MaE recognition, age estimation, and face attribute prediction. Similarly, Foggia et al. (Foggia et al., 2023) proposed a comprehensive multi-task framework to integrate gender, age, ethnicity, and MaE recognition using facial images.
@@ 4.1.4. Attention-based learning
Attention-based learning is used to enrich the spatial information in deep static MaE recognition. The attention modules can focus on both global and local regions and generate more comprehensive static facial expression features that typically include local information, global information, and the corresponding dependency. Drawing inspiration from the human ability to locate salient objects in complex visual environments, attention mechanisms have emerged as a powerful tool in deep static MaE recognition (Itti and Koch, 2001). By prioritizing critical facial regions, attention modules can significantly enhance the extraction and integration of expression features in order to achieve a more accurate and robust recognition model. For example, Fernandez et al. (Marrero Fernandez et al., 2019) developed an attention module to enhance the extraction of expression features by assigning higher weights to relevant regions. Li et al. (Li et al., 2020a) integrated LBP and deep features with an attention mechanism to enhance the performance of deep static MaE recognition. In wild scenarios, pose variations and occlusions present significant challenges for deep static MaE recognition (Li et al., 2019c; Wang et al., 2020b). To address the aforementioned issues, various attention modules have been proposed to enhance model robustness and performance. For example, Li et al. (Li et al., 2019c) introduced an attention module designed to identify local facial regions associated with MaEs while simultaneously incorporating complementary global facial information. This dual-focus approach enables more robust MaE recognition under occlusion conditions by balancing localized details with a holistic view of the face. Wang et al. (Wang et al., 2020b) proposed a region attention module that is tailored to address occlusions and pose variations. The region attention module in (Wang et al., 2020b) highlights significant facial regions by leveraging attention weights and combines local region features to generate a compact and fixed-length feature representation for final classification. Zhao et al. (Zhao et al., 2021c) introduce a global–local attention module with multi‑scale receptive fields. The global branch gathers broad contextual cues, while the local branch highlights the most salient facial regions. Together, these two streams yield a richer and more discriminative representation for facial‑expression recognition. Liu et al. (Liu et al., 2022a) design an adaptive, multi‑layer perceptual attention module that weights facial attributes according to human‑perception cues. By jointly emphasizing global context and salient local details, the module markedly improves the robustness of static MaE recognition.
@@
First proposed in (Dosovitskiy et al., 2020), vision transformers (ViTs) have brought remarkable advancements in the domain of deep static MaE recognition, primarily due to their ability to capture long-range dependencies across image patches (Aouayeb et al., 2021; Zheng et al., 2023). By leveraging the self-attention module inherent in transformers, Aouayeb et al. (Aouayeb et al., 2021) pioneered the integration of ViTs with a squeeze-and-excitation block that can extract expression-related features. Xue et al. (Xue et al., 2021) exploited the ViT architecture to explore relationships among various facial regions for a more holistic understanding of static MaEs. To address the occlusion issue, Liu et al. (Liu et al., 2023a) introduced a patch attention module that assigns attention weights to local facial regions. The integration of the patch attention module with a ViT can effectively capture both local and global dependencies. Building upon the advancements of ViTs, Zheng et al. (Zheng et al., 2023) developed a dual-stream model that leverages a cross-fusion mechanism to combine facial landmark features with image-based features. Furthermore, the incorporation of additional modalities has proven instrumental in augmenting transformer-based deep static MaE recognition. For example, gradient images (Meng and Shi, 2024) and gray-level co-occurrence matrices (Sun et al., 2023b) have been used as complementary inputs to enrich the feature space. These modalities not only enhance the discriminative power of the models but also provide deeper insights into the complex interplay of facial features. The optimization of transformer architectures also plays a critical role in advancing deep static MaE recognition. In (Feng et al., 2023), Feng et al. demonstrated that adjusting the model structures of ViTs through optimization algorithms can induce significant performance improvements.
@@ 4.1.5. Self-supervised learning
The self-supervised learning has also been applied to multi-modal MaE recognition to address the challenges of integrating data from various modalities without explicit annotation. Siriwardhana et al. (Siriwardhana et al., 2020) utilized pre-trained self-supervised models to extract text, speech, and vision features to improve deep static MaE recognition tasks. Moreover, multi-modal self-supervised learning frameworks have been explored to extract and fuse multi-modal features to achieve promising results in deep static MaE recognition (Chaudhari et al., 2023; Halawa et al., 2024).
4.1.6. Integrated Comparisons & Practical Insights on Deep Static MaE Recognition
The transfer learning, multi-task learning, and self-supervised learning paradigms can enhance the deep static MaE recognition by feature generalization. While enriching feature representation via pre-training (transfer learning (Ng et al., 2015; Deng et al., 2009; Akhand et al., 2021; Ngo and Yoon, 2020; Atabansi et al., 2021; Cao et al., 2018)), multi-objective optimization (multi-task learning (Chen et al., 2021; Pons and Masip, 2018b; Yu and Xu, 2022; Li et al., 2023; Yu et al., 2020; Xiao et al., 2023; Zhao et al., 2021a; Chen et al., 2022a; Zaghbani and Bouhlel, 2022; Pan et al., 2022; Qin et al., 2024; Foggia et al., 2023)), and unlabeled data exploitation (self-supervised learning (Li et al., 2022c; Wang et al., 2022a; Chen et al., 2023b; Fang et al., 2023; An and Jia, 2024; Roy and Etemad, 2021, 2023; Zbontar et al., 2021; Siriwardhana et al., 2020; Chaudhari et al., 2023; Halawa et al., 2024)), the aforementioned learning paradigms are still hindered by domain shifts, insufficient labeled data, and intricate training processes. Complementary techniques (e.g., ensemble learning (Kim et al., 2016; Sharifnejad et al., 2021; Hariri and Farah, 2021; Mohan et al., 2021; Shao and Qian, 2019; Viswanatha Reddy et al., 2020; Kim et al., 2015; Pons and Masip, 2018a; Xie and Hu, 2019; Wadhawan and Gandhi, 2023; Hua et al., 2019; Saurav et al., 2022) and attention-based learning (Itti and Koch, 2001; Marrero Fernandez et al., 2019; Li et al., 2020a, 2019c; Wang et al., 2020b; Zhao et al., 2021c; Liu et al., 2022a; Dosovitskiy et al., 2020; Aouayeb et al., 2021; Zheng et al., 2023; Xue et al., 2021; Liu et al., 2023a; Meng and Shi, 2024; Sun et al., 2023b; Feng et al., 2023)) can further improve accuracy by enhancing feature representation and contextual understanding. However, these methods often come with substantial computational costs due to the use of multiple backbones, additional processing pipelines, and more complex model architectures. For a detailed comparison of the strengths and weaknesses, please refer to Appendix 3.1.
A central challenge in deep static MaE recognition lies at balancing accuracy with computational efficiency. Complex models usually yield good accuracy. For example, transformer‑based learning can achieve state‑of‑the‑art results by modeling complicated context information (Aouayeb et al., 2021; Zheng et al., 2023; Xue et al., 2021). However, the substantial computing loads are burdensome for the low-power IoT devices. Two major strategies have been explored, i.e., model simplification and computation offloading. For the model simplification, lightweight and pruned networks preserve most of the accuracy while sharply reducing parameters and computation (Shao and Qian, 2019; Wadhawan and Gandhi, 2023). For the computing offloading, inference is (partially) wholly migrated to edge or cloud servers with higher compute capability, and thereby reduce the local latency without sacrificing performance (Wadhawan and Gandhi, 2023). A second major hurdle is reliable MaE recognition in the wild, where frequent occlusions and pose variations present. One effective solution is to construct robust feature representations to occlusions and pose variations. This can be achieved through multi-task learning (Chen et al., 2022a; Zaghbani and Bouhlel, 2022, 2022) and self-supervised learning (Roy and Etemad, 2021, 2023; Zbontar et al., 2021), which help the model capture robust, task-related features. Additionally, attention-based learning (Li et al., 2019c; Wang et al., 2020b) can be employed to enhance occlusion awareness by computing attention weights across global and local features, allowing the model to dynamically focus on unoccluded and informative regions of the face. These strategies push static MaE recognition toward robust, resource‑aware deployment in the wild.
@@ 4.1.6. Integrated Comparisons & Practical Insights on Deep Static MaE Recognition
Explicit Spatio-temporal Learning Zhang et al. (Zhang et al., 2017) 2017 DA N/A CK+/Oulu-CASIA/MMI 98.5%/86.25%/81.18% CNN+RNN 10 fold
@@
Hasani et al. (Hasani and Mahoor, 2017) 2017 FA+DA RES CK+/MMI/FERA/DISFA 93.21%/77.50%/77.42%/58.00% CNN+LSTM 5 fold
@@
Multi-task Learning Yu et al. (Yu et al., 2020) 2020 FA+FD N/A CK+/Oulu-CASIA 98.77%/90.40% CNN+LSTM 10 fold
Jin et al. (Jin et al., 2021) 2021 DA RES Aff-wild2 77.09% CNN Hold out
Xie et al. (Xie et al., 2023) 2023 FA+DA+ROI Trans+Atten CMU-MOSI/CMU-MOSEI 85.36%/84.61% BERT Hold out
Attention-Based Learning Meng et al. (Meng et al., 2019) 2019 FD+FA RES+Atten CK+/AFEW 99.69%/51.18% CNN 10 fold
@@
Huang, et al. (Huang et al., 2021) 2021 FA RES+Trans CK+/FERplus/RAF-DB 100%/90.04%/88.26% CNN 10 fold+Hold out
@@
Chen et al. (Chen et al., 2024) 2024 DA RES+Trans DFEW/FERV39k/AFEW
@@
Zhang et al. (Zhang et al., 2024) 2023 FD+DA Trans MAFW11/MAFW43/DFEW/AFEW
@@
Sun et al. (Sun et al., 2023a) 2023 ROI Trans DFEW/FERV39k/MAFW
@@
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; FA: Face alignment; N/A: Not applicable; UAR: Unweighted average recall; WAR: Weighted average recall; RMSE: Root mean squared error; CC: Correlation coefficient; CCC: Concordance correlation coefficient.
  • 2Trans: Transformer; Atten: Attention; nsCNN: n-stream CNN; LOSO: Leave-one-subject-out; RNN: Recurrent neural network; LSTM: Long short-term memory; BERT: Bidirectional encoder representations from transformer; GCN: Graph convolutional network.